Royal Lawncare Company produces and sells two packaged products—Weedban and Greengrow. Revenue and cost information relating to the products follow:

Product

Weedban Greengrow
Selling price per unit $ 12.00 $ 37.00
Variable expenses per unit $ 2.80 $ 11.00
Traceable fixed expenses per year $ 136,000 $ 43,000

Common fixed expenses in the company total $99,000 annually. Last year the company produced and sold 42,000 units of Weedban and 16,000 units of Greengrow.

Required:

Prepare a contribution format income statement segmented by product lines.

Product Line
Total Company Weedban Greengrow
Answer #1

Solution

Income Statement
Product Line
Details Total Company Weedband Greengrow
Sales $ 1,096,000 $ 504,000 $ 592,000
Variable Cost $ 293,600 $ 117,600 $ 176,000
Contribution margin $ 802,400 $ 386,400 $ 416,000
Traceable Fixed cost $ 179,000 $ 136,000 $ 43,000
Income Before Common Fixed Cost $ 623,400 $ 250,400 $ 373,000
Common Fixed Cost $ 99,000
Net income $ 524,400

Working

Income Statement
Product Line
Details Weedband Greengrow
Sales =42000*12 =16000*37
Variable Cost =42000*2.8 =16000*11